# Towards Coordinate- and Dimension-Agnostic Machine Learning for Partial Differential Equations

### v1

@@ 3.1 The FitzHugh-Nagumo System
In the FitzHugh-Nagumo system bar2003bifurcation two scalar fields participate: the excitation variable $V(x,t)$ and the recovery variable $W(x,t)$. The spatiotemporal dynamics of these fields are governed by the set of two coupled PDEs:
@@
where we set $(\varepsilon,\beta,\gamma)=(0.02,0,0.5)$. For this choice of parameters, the system generates spiral waves in two-dimensional space and scroll waves in three-dimensional space, whereas such phenomena do not emerge in dimension one. In coordinate-free form, the PDE can be expressed as:
where the location $x$ can be omitted from the equation since this is a spatial coordinate-free formalism.
For brevity, we express all PDEs from this point onward in coordinate-free representation.
3.2 The Barkley System
The Barkley reaction-diffusion system barkley1991model models excitable media and oscillatory media. It is often used as a qualitative model in pattern forming systems like the Belousov-Zhabotinsky reaction and other systems that are well described by the interaction of an activator and an inhibitor component. In the simplest case, the Barkley system is described by a set of two coupled PDEs:
where we choose the parameters $(\varepsilon,a,b)=(0.02,0.75,0.02)$ so that this system can generate spiral and scroll waves in two- and three-dimensional spaces respectively. For a coordinate-explicit form of the equation, see Appendix C.
3.3 A Modified Patlak-Keller-Segel System
Recent technological advancements have enabled us to directly measure in situ the time-evolution dynamics of bacterial density and the concentration of environmental cues to which the bacteria respond, with high precision phan2024direct . We can therefore develop much more accurate, macroscopic mathematical models to describe the collective behavior of bacterial swarms. We consider the modified Patlak-Keller-Segel model for bacterial chemotaxis as proposed in phan2024direct , supported by direct in situ observations of E. coli (Escherichia coli) density and Asp (Aspartate) chemoattractants. The dynamics are described by a set of two PDEs, which can be can be written in a non-dimensionalized form as follows:
@@ 3.3 A Modified Patlak-Keller-Segel System
For a coordinate-explicit form of the equation, see Appendix C.
4 Computational Experiments
For each of the systems of Section 3 we follow the same workflow:
@@ 4 Computational Experiments
    Given training data of the scalar fields $\quantity{\Psi_{i}(x,t)}_{i=0}^{N}$ in one-dimensional Euclidean space, we project the data onto the spatial coordinate-free basis $\mathcal{B}$ of (5) and estimate the time derivatives $\{\partial\hat{\Psi}_{i}/\partial t\}$ using a finite-difference scheme,
@@
    We train a fully-connected neural network ^$\hat{\mathcal{F}}_{\text{NN}}$ to predict $\{\partial\hat{\Psi}_{i}/\partial t\}$ from the basis $\mathcal{B}$,
@@
    Given the trained network ^$\hat{\mathcal{F}}_{\text{NN}}$, we integrate the corresponding scalar fields (now defined on different choices of domain) given initial conditions.
We provide numerical examples for the following domains:
@@
    Section 4.1: 1-Dimensional Euclidean Space (periodic boundary conditions) with Cartesian and non-Cartesian coordinates,
@@
    Section 4.2: 2-Dimensional Euclidean Space (periodic boundary conditions) with Cartesian and non-Cartesian coordinates, and 3-Dimensional Euclidean Space (Neumann no-flux boundary conditions) with Cartesian coordinates,
@@
    Section 4.3: 2-Dimensional intrinsically curved surface (heterogeneous boundary conditions) embedded in 3-Dimensional Euclidean Space with geographic and stereographic coordinates,
The learning of all PDE systems is conducted on data from a one-dimensional periodic domain $x\in[-L,+L]$. The data and neural network architectures are detailed in Appendix D. All simulations use evenly spaced grids for finite-difference computations in the specified coordinate system, and we verify that the numerical schemes are robust to space-time discretizations (see Appendix H). For notational convenience, we denote the fields obtained by numerically integrating the exact PDE as $\Psi_{i}$, and the predictions generated by the neural network models as $\hat{\Psi}_{i}$. The argument $\eta(\zeta)$ denotes that the computation was performed in $\zeta$-coordinates and has been push-forwarded into $\eta$-coordinates via a coordinate transformation. A uniform grid in $\zeta$-coordinates will in general correspond to a non-uniform grid in $\eta$-coordinates.
All quantitative comparisons between the “true” and learned dynamics are contained in Appendix J.
4.1 One-Dimensional Euclidean Space
$x=f(y)\equiv x(y)=(y/2)\left[1+(y/\lambda)^{2}\right]\ ,$ (11)
as illustrated in Fig.2A. We set $\lambda=L$. The inverse function $y=f^{-1}(x)\equiv y(x)$ is provided in Appendix F.1. The input pre-processing for the neural network (Fig. 1B, in $y$-coordinate) is detailed in Appendix F.2.
We test the trained model on a new initial condition $t=0$ with uniform $A$ and localized peaks in $B$ (see Appendix H.1). The true dynamics, $A(x,t)$ and $B(x,t)$, are obtained by integrating Eq.(29) (Fig.2B1); the learned dynamics, $\hat{A}(x,t)$ and $\hat{B}(x,t)$, are generated by the neural network (Fig.2B2) are a good approximation. To test our coordinate-free generalization claims, we interpolate the initial field configurations to $y$-coordinates (see Appendix G), integrate the neural network model to get $\hat{A}(y,t)$ and $\hat{B}(y,t)$ (Fig.2D1), and then map the results back to $x$-space to arrive at $\hat{A}(x(y),t)$ and $\hat{B}(x(y),t)$ (Fig.2D2). We can also transform Fig. 1B1 to $y$-coordinates, obtaining $A(y(x),t)$ and $B(y(x),t)$ as shown in Fig. 2C. We observe that Fig. 2C and Fig. 2D1 agree, as well as Fig. 2A1 and Fig. 2D2, demonstrating that our machine learning method generalizes across coordinate systems.
4.2 Two- and Three-Dimensional Euclidean Space
$x_{1}=f(y_{1})\equiv x_{1}(y_{1})\ ,\ x_{2}=f(y_{2})\equiv x_{2}(y_{2})\ ,$ (12)
where $f$ is as in Eq.(11) and we set $\lambda=L$. The domain is the square $(x_{1},x_{2})\in[-L,+L]\times[-L,+L]$ with periodic boundaries. An illustration of $(y_{1},y_{2})$ versus $(x_{1},x_{2})$ is in Appendix F.1; input pre-processing in $(y_{1},y_{2})$-coordinates for the neural network is detailed in Appendix F.3.
Spiral waves can emerge for certain instances of the FitzHugh-Nagumo system. We choose an initial field configuration at $t=0$ where the spiral wave pattern is clearly visible (an explanation of how we do it is provided in Appendix H.2). In $(x_{1},x_{2})$-coordinates, the true dynamics are computed by integrating Eq.(6) (Fig. 3A1-4 top row, $V(x_{1},x_{2},t)$ and $W(x_{1},x_{2},t)$). In $(y_{1},y_{2})$-coordinates, the learned model is integrated from the extrapolated initial condition (see Appendix G for details), and the results are mapped back to $(x_{1},x_{2})$ (Fig. 3A1-4 bottom row, $\hat{V}(x_{1}(y_{1}),x_{2}(y_{2}),t)$ and $\hat{W}(x_{1}(y_{1}),x_{2}(y_{2}),t)$). We observe that the predictions agree, suggesting that our machine learning method not only generalizes across coordinate systems, but also it extends across dimensions. The correct predictions are made in two dimensions, even though the model was trained in one dimension, indicating that the approach is dimension-free.
$\star d\Psi_{i}\big{|}_{\partial\mathcal{C}}=0\ ,$ (13)
We initialize the Barkley system with a field configuration that clearly displays scroll wave patterns (Appendix H.3). Fig. 3B1–4 shows that the neural network predictions, $\hat{U}(x_{1}(y_{1}),x_{2}(y_{2}),x_{3}(y_{3}),t)$ and $\hat{V}(x_{1}(y_{1}),x_{2}(y_{2}),x_{3}(y_{3}),t)$, closely match the true dynamics, $U(x_{1},x_{2},x_{3},t)$ and $V(x_{1},x_{2},x_{3},t)$, as governed by Eq. (28), demonstrating the adaptability of our dimension-free model to diverse boundary conditions.
4.3 Two-Dimensional Curved Surface Embedded in Three-Dimensional Euclidean Space
Finally, in addition to increasing the number of the spatial dimensions and modifying the boundary conditions, we now bend the space (intrinsically). For this case, we explore the dynamics of the FitzHugh–Nagumo fields on a two-dimensional domain (similar to Fig. 3A but not flat) embedded in a spherical surface — a manifold with constant positive intrinsic curvature. We consider two coordinate representations of this curved space — the geographic $(\theta,\phi)$ and the stereographic $(\rho,\varphi)$ systems — as illustrated in Fig. 4A-C (See Appendix I for more details). The transformation between these coordinate systems is given by:
$\rho=R\tan\left(\phi/2\right)\ ,\ \varphi=\phi\ ,$ (14)
The spatial domain will be constrained between the radial values $\rho_{\min}$ and $\rho_{\max}$, and the angular values $\varphi_{\min}$ and $\varphi_{\max}$, which can be determined from $\theta_{\min}$, $\theta_{\max}$, $\phi_{\min}$, $\phi_{\max}$ (Appendix H.4), We impose heterogeneous boundary conditions: the Neumann (no-flux) conditions on the latitude boundaries and the periodic conditions on the longitude boundaries.
We initialize the system with a clearly defined spiral pattern for the fields $V$ and $W$ (defined in $(\theta,\phi)$-coordinates, see Appendix H.4). The true dynamics — $V(x(\theta,\phi),y(\theta,\phi),z(\theta,\phi),t)$ and $W(x(\theta,\phi),y(\theta,\phi),z(\theta,\phi),t)$ — are obtained by integrating Eq.(6) in $(\theta,\phi)$-coordinates and then mapped to the embedding space $(x,y,z)$, as illustrated in the top row of Fig.4D1–4. For the stereographic $(\rho,\varphi)$-coordinates, we extrapolate the initial condition (Appendix G), integrate the neural network learned dynamics, and transform the results to $(x,y,z)$ to obtain $\hat{V}(x(\rho,\varphi),y(\rho,\varphi),z(\rho,\varphi),t)$ and $\hat{W}(x(\rho,\varphi),y(\rho,\varphi),z(\rho,\varphi),t)$, shown in the bottom row of Fig.4D1–4. We observe a remarkable agreement (quantified in Appendix J) between both predictions, illustrating that our machine learning method not only generalizes across coordinate systems, extends across dimensions, and adapts to different boundary conditions, but also remains robust under curvatures.
@@ 5 Discussion
We have shown that, for certain PDE systems, our approach enables a neural network to learn the evolution operator in a “spatially liberated” manner. The training can be performed using data from simple 1-D numerical simulations or experimental settings, yet it can generalize across dimensions, coordinate systems, and domain curvatures, even when varying boundary conditions. Training can also be performed “fusing" available data from multiple dimensions or geometries. This flexibility makes our approach particularly valuable at the interface between experiments and engineering, as it enables seamless integration of experimental data from different geometry/dimension experiments into predictive models, allowing theoretical frameworks to adapt to complex real-world geometries.
Here, we only focus on second-order PDEs of scalar fields that are invariant under the full local symmetry group of Euclidean space – that is, local orthogonal transformations (rotations and reflections) and local translations (interpreted in the diffeomorphic sense) — which often emerge from long-wavelength physical effective field theories kaplan2005five ; buchalla2014power . Next, it would be interesting to consider PDE systems beyond these limitations, such as those involving vector or tensor fields, where additional symmetries and gauge constraints play a role. Further directions include investigating systems with higher-order derivatives, nonlocal interactions, or anisotropic effects to assess the broader applicability of spatially liberated machine learning approach. Notably, we have only considered integer dimensions $n\in\mathbb{N}$, but it would be worthwhile to explore generalizations to fractional dimensions karniadakis2019handbook ; lischke2018fractional ; kochubei2019handbook ; phan2024vanishing . Additionally, real-world experiments often involve mixed boundary conditions, such as Robin-type boundary conditions netuka1971robin ; daners2013principal ; papageorgiou2014multiple ; d2016multiple ; robin ; phan2024social , which could present another avenue for future investigation. Applications to elliptic, hyperbolic, and more general “mixed” PDE operators can and should also be pursued.
Acknowledgement
The work of I.G.K. and T.V.P. was partially supported by the U.S. Department of Energy and the U.S. National Science Foundation. S.V. was partly supported by NSF CAREER 2339682 and NSF BSF 2430292. We wish to acknowledge valuable discussions with Prof. Dwight Barkley (University of Warwick), who kindly provided us with code for the Barkley model. This preprint has been assigned the number LA-UR-25-24791.
@@ Appendix A Background on Differential Forms
The Hodge star is the operator that turns a $k$-form into a $(n-k)$-form:
@@
In a coordinate-explicit representation, given the metric structure $g$ defined on the manifold $\mathcal{M}$, the Laplacians and the inner products can be written as:
@@
    $\star\ d\star d\Psi_{j}=\left|g\right|^{-1/2}\partial_{\mu}\left(\left|g\right|^{1/2}g^{\mu\nu}\partial_{\nu}\Psi_{j}\right)\ ,$ (21)
@@
  • The inner products:
    $\langle d\Psi_{j},d\Psi_{k}\rangle\equiv\star\left(d\Psi_{j}\wedge\star\ d\Psi_{k}\right)=g^{\mu\nu}\partial_{\mu}\Psi_{j}\partial_{\nu}\Psi_{k}\ ,$ (22)

### v2

@@ 3.1 Time-Evolution Systems
The FitzHugh-Nagumo System. In the FitzHugh-Nagumo system (Bär et al., 2003) two scalar fields participate: the excitation variable $V(x,t)$ and the recovery variable $W(x,t)$. The spatiotemporal dynamics of these fields are governed by the set of two coupled PDEs:
@@
where we set $(\varepsilon,\beta,\gamma)=(0.02,0,0.5)$. For this choice of parameters, the system generates spiral waves in two-dimensional space and scroll waves in three-dimensional space, whereas such phenomena do not emerge in dimension one.
The Barkley System. The Barkley reaction-diffusion system (Barkley, 1991) models excitable media and oscillatory media. It is often used as a qualitative model in pattern forming systems like the Belousov-Zhabotinsky reaction and other systems that are well described by the interaction of an activator and an inhibitor component. In the simplest case, the Barkley system is described by a set of two coupled PDEs:
where we choose the parameters $(\varepsilon,a,b)=(0.02,0.75,0.02)$ so that this system can generate spiral and scroll waves in two- and three-dimensional spaces respectively.
A Modified Patlak-Keller-Segel System. Recent technological advancements have enabled us to directly measure in situ the time-evolution dynamics of bacterial density and the concentration of environmental cues to which the bacteria respond, with high precision (Phan et al., 2024). We can therefore develop much more accurate macroscopic mathematical models to describe the collective behavior of bacterial swarms. We consider the modified Patlak-Keller-Segel model for bacterial chemotaxis as proposed in Phan et al. (2024), supported by direct in situ observations of E. coli (Escherichia coli) density and Asp (Aspartate) chemoattractants. The dynamics are described by a set of two PDEs, which can be can be written in a non-dimensionalized form as follows:
@@
Method. To study all these systems, we use the following algorithmic workflow:
@@
    Given training data of the scalar fields $\quantity{\Psi_{i}(x,t)}_{i=0}^{N}$ in one-dimensional Euclidean space, we project the data onto the spatial coordinate-free basis $\mathcal{B}$ in (1) and estimate the time derivatives $\{\partial\hat{\Psi}_{i}/\partial t\}$ using a finite-difference scheme,
@@
    We train a fully-connected neural network $\hat{\mathcal{F}}_{\text{NN}}$ to predict $\{\partial\hat{\Psi}_{i}/\partial t\}$ from the basis $\mathcal{B}$,
@@
    Given the trained network $\hat{\mathcal{F}}_{\text{NN}}$, we integrate the corresponding scalar fields (now defined on different choices of domain) given initial conditions.
One-Dimensional Euclidean Space. We first study the generalization of our coordinate-free representation to arbitrary (nonlinear) parametrizations in the spacial domain. In Figure 2, we demonstrate that the learned representation can be used to accurately predict the time evolution for the Patlak-Keller-Segel system under different choices of coordinates. For details, see Appendix D.1.
Two- and Three- Dimensional Euclidean Space. We study the generalization of the learned FitzHugh-Nagumo and Barkley systems in two and three dimensions, represented in Figure 3. Again, we observe agreement between the true and generalized (learned from 1-D) dynamics. We emphasize that the learned operator produces spiral waves, which are a property of solutions in higher dimensions that cannot be inferred from the 1-D solution. For details, see Appendix D.2.
Generalization to different curvatures and topologies. The proposed coordinate-free representation generalizes to domains with different curvature and topology. In particular, we demonstrate that the learned (from 1-D data) operator accurately predicts the time-evolution of the FitzHugh-Nagumo system on a 2-D domain embedded on the surface of a 3-D ball with nontrivial curvature (see Appendix D.3, Figure 6). In this example, we consider both the geographic and stereographic coordinates of the sphere. In addition, we also consider the case where the domain is a flat punctured disk with nontrivial topology. For details, see Appendix D.4.
3.2 General PDE Systems
The Helmholtz Equation. Given data satisfying the Helmholtz equation $\Delta u=-ku$, for $k=1$, the goal is to learn the PDE from data in dimension 1 (where it reduces to the harmonic oscillator), and generalize to higher dimensions. The coordinate-free features for one scalar field are $\mathcal{B}=\quantity{u,\expectationvalue{\gradient u,\gradient u},\Delta u}$. Since we only have one PDE, $\mathcal{B}(u)$ should span a two-dimensional manifold that we call $\mathcal{L}$, and since the PDE is linear, PCA should be able to identify it. At this point, the operator is implicitly encoded in the weights and components of the fitted PCA model. (Also note that, after computing derivatives, and embedding in $\mathcal{B}$, we discard all information that has to do with which point corresponds to which solution; they all simply lie on a submanifold in that space.)
Now, suppose that we want to transfer this operator to another domain, for example a 2D unit square equipped with the flat metric. We will do this using a Physics-Informed neural network (PINN) (Raissi et al., 2019), solving the PDE implicitly defined by constraints in the manifold $\mathcal{L}$ and boundary conditions prescribed by the user. In the classical setting, for any given domain with an initial condition, the PINN produces a candidate solution $\hat{u}$. The PDE law and boundary conditions are imposed in the loss function as ‘soft’ constraints. To train the PINN using the implicit coordinate-free representation of the operator, encoded by the prior PCA components, we proceed as follows:
@@ 3.2 General PDE Systems
    Sample a set of points $\quantity{x}$ in the domain $\Omega$
@@
[…]
@@
[…]
  4. 4.
[…]
  5. 5.
[…]
    $\mathcal{E}_{\text{PDE}}=\norm{\hat{\hat{u}}_{\mathcal{B}}-\mathcal{B}(\hat{u})}$.
The proposed framework is depicted in Fig. 4. The boundary conditions (BC) are imposed in exactly the same manner as in the classical setting, with a PINN training loss taking the form $\mathcal{E}_{\text{PINN}}=\mathcal{E}_{\text{PDE}}+\mathcal{E}_{\text{BC}}$.
In Fig. 14 (Appendix M) we compare solutions to the 2D Helmholtz equation on the unit square with the flat metric, obtained using (i) PINNs and (ii) the projection-based algorithm described above, where the PDE law is implicitly encoded by the fixed PCA model in coordinate-free space. For details, see Appendix M.
Figure 4: Implicit Coordinate-Free Representation: Left: Solutions of the 1-D Helmloltz problem (harmonic oscillator) in the original variables. Middle: Embedding of the solutions to $\mathcal{B}$, with black arrows depicting the two PCA modes obtained from the data. Right: Embedding of one prediction of a randomly initialized PINN in $\mathcal{B}$ (yellow), along with its projection (red) on the previously learned PCA submanifold . The PDE loss tests whether the yellow and green representations agree, i.e. if the PINN solution resides on the learned submanifold.
Remark: While the PDE constraint itself can be encoded as a relation (linear or nonlinear) between the coordinate-free basis elements in any input space, it does not inform us about what the necessary and sufficient initial and boundary conditions are to solve a PDE in a different geometry or dimension. For a relevant discussion in terms of well-posedness, see Bertalan et al. (2025)). This is not the case for time-evolution problems, where knowledge of the initial field is sufficient.
3.3 Vector PDEs
Figure 5: Magnitude of velocity field for simulated flow around cylinder (top) and integrated (learned, coordinate-free) Taylor-Green Vortex dynamics (bottom)
Incompressible Navier Stokes: We apply our extended methodology to the 2D incompressible Navier-Stokes equations. We consider two well-studied settings, that of flow around a cylinder, and the Taylor-Green Vortex (Taylor & Green, 1937), illustrated in Figure 5. In the latter case, an analytical solution is known. The equation takes the form:
$\partialderivative{v}{t}=-\quantity(v\cdot\gradient)v-\frac{1}{\rho}\gradient p+\frac{\eta}{\rho}\Delta v$ (8)
$\gradient\cdot{v}=0,$ (9)
where $v=(v_{1},v_{2})$ is the velocity vector field, $p$ is the pressure, a scalar quantity, $\rho$ is the density which is assumed to be constant, and $\eta$ is the dynamic coefficient of the incompressible fluid, and is also constant. It is easy to see that this is an instance of the general form (3).
We will consider the problem of learning the momentum equation from the flow around a cylinder, and transferring it to the taylor-green vortex setting. Because of the simplified nature of the incompressible equations, we can perform least-squares to identify the coefficients of the vector features (which play the role of the $f_{i}$ in (3)) instead of using a neural network. We note that the compressible Navier-Stokes would require the $f_{i}$ be nonlinear, at least as a function of density. We follow a familiar pattern: We compute the coordinate-free features of the Navier-Stokes solution in the cylinder setting, and perform least-squares to learn $\partialderivative{v}{t}$ as a function of the spatial-derivative features. We then transfer the operator to the Taylor-Green Vortex setting (assuming the pressure is known), and integrate given an initial condition for which the analytical solution is known. In these two examples, the topology of the domain is different. For details, see Appendix N.
@@ 4 Discussion
We propose a framework that, for a general class of PDE systems, allows us to learn and represent the equation in a “spatially liberated” manner. The training can be performed using data from simple 1-D numerical simulations or experimental settings, yet it can generalize across dimensions, coordinate systems, and domain curvatures, even when varying boundary conditions. Training can also be performed “fusing" available data from multiple dimensions or geometries. This flexibility makes our approach particularly valuable at the interface between experiments and engineering, as it enables seamless integration of experimental data from different geometry/dimension experiments into predictive models, allowing theoretical frameworks to adapt to complex real-world geometries.
Using ideas from any-dimensional machine learning, in this work, we focus on identification of second-order PDEs that are invariant under the full local symmetry group of Euclidean space (rotations and reflections, and translations) – which often emerge from long-wavelength physical effective field theories (Kaplan, 2005; Buchalla et al., 2014).The next step is to apply this framework to solving PDEs in a dimension agnostic manner that do not explicitly require the use of integrators (here we use coordinate-free integrators). For example, by defining an any-dimensional formulation of neural operators or PINNs that allow us to transfer PDE solutions from one domain to another.
Implicit to any-dimensional generalization is the underlying assumption of what it means to take a model from one space to another. In the case of graph neural networks assumptions that GNNs make to generalize to different graph sizes have been studied extensively in the graph transferability literature (Ruiz et al., 2020; Maskey et al., 2023; Velasco et al., 2024; Levin et al., 2025). For PDEs, the generalization to other domains is done via the intrinsic manifold operators (Laplacians, exterior derivatives, etc), but more work is needed to fully characterize what these methods are doing.
We expect this type of generalization to be broadly beneficial for the machine learning community, as PDEs play a central role not only in modeling of physical systems, but also optimization algorithms, sampling, and diffusion algorithms, which are fundamental to generative AI. A theoretically grounded approach to distilling information that can be applied to systems of arbitrary dimension could significantly boost efficiency for high-dimensional methods.
@@ Appendix A Background on Differential Forms
The natural mathematical formalism to express coordinate-free operators on Riemannian manifolds is given by exterior calculus, which is built on differential forms. In this section, we give the relevant definitions and background.
@@
The Hodge star is the operator that turns a $k$-form into a (nk)(n-k)-form:
@@
In a coordinate-explicit representation, given the metric structure $g$ defined on the manifold $\mathcal{M}$, the Laplacians and the inner products of gradients from Section 2 can be written as:
@@
    $\Delta\Psi_{j}=\star\ d\star d\Psi_{j}=\left|g\right|^{-1/2}\partial_{\mu}\left(\left|g\right|^{1/2}g^{\mu\nu}\partial_{\nu}\Psi_{j}\right)\ ,$ (15)
@@
  • The inner products of the gradients:
    $\langle\nabla\Psi_{j},\nabla\Psi_{k}\rangle=\langle d\Psi_{j},d\Psi_{k}\rangle\equiv\star\left(d\Psi_{j}\wedge\star\ d\Psi_{k}\right)=g^{\mu\nu}\partial_{\mu}\Psi_{j}\partial_{\nu}\Psi_{k}\ ,$ (16)